# Fractional binomial distributions induced by the generalized binomial theorem and their applications

### v1

@@ 2.3. Uniform bound of $Z_{\alpha,x}^{(n)}$
If $\alpha\in(0,1)$, then Proposition 1.1 implies that $Z_{\alpha,x}^{(n)}\leq 1$ for $x\in[0,1]$ and $n\in\mathbb{N}$. On the other hand, Equation (1.7) in [GS23, Theorem 1.3] leads to $Z_{\alpha,x}^{(n)}\leq\alpha$ for $x\in[0,1]$ and $n\in\mathbb{N}$ when $\alpha\geq 1$. Hence, the uniform upper bounds of (2.25) and (2.26) have been obtained.
We next discuss the lower bounds of $Z_{\alpha,x}^{(n)}$. Because it holds that $Z_{\alpha,x}^{(n)}=Z_{\alpha,1-x}^{(n)}$ and $Z_{\alpha,0}^{(n)}=\alpha$, it suffices to consider only the case where $x\in(0,1/2]$. It follows from [HH10, Proposition 2.6] that
@@
$-\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}+1\in\begin{cases}\left(0,(1-\alpha)(1-x)^{\alpha n}\right)\subset(0,1-\alpha)&\text{if }\alpha\in(0,1),\\ \left((1-\alpha)(1-x)^{\alpha n},0\right)&\text{if }\alpha\in(1,2).\end{cases}$
@@
$\geq\alpha\left(1-\frac{1}{2}\right)^{2\alpha nx}=\alpha\times 2^{-2\alpha nx},\qquad x\in(0,1/2],\,n\in\mathbb{N}$ (2.28)
@@
$Z_{\alpha,x}^{(n)}\geq\alpha\times 2^{-2\gamma},\qquad x\in(0,1/2],\,n\in\mathbb{N}\,\,\text{with $\alpha nx\leq\gamma$.}$
We now deal with the upper bound of $\sum_{\omega\in K_{\alpha}\setminus\{1\}}(1-x+x\omega)^{\alpha n}$ for $x\in(0,1/2]$ and $n\in\mathbb{N}$ with $\alpha nx>\gamma$. From (2.18), we have
@@
for $x\in(0,1/2]$ and $n\in\mathbb{N}$, where we used #$\#(K_{\alpha})=\lfloor\alpha/2\rfloor+\lceil\alpha/2\rceil\leq\alpha+1$ in the fourth line and
@@
for $x\in(0,1/2]$ and $n\in\mathbb{N}$. ∎
2.4. Uniform estimates for moments of the fractional binomial distribution
@@
$\hskip 28.45274pt-(\alpha n)^{-m}\sum_{k=0}^{m}W_{-\alpha nx}(m,k)(\alpha n)_{k}x^{k}\Bigg{|}=O\left(\frac{1}{n^{m}}\right)$ (2.30)
@@
$\hskip 28.45274pt+\sum_{k=0}^{m}W_{-\alpha nx}(m,k)(\alpha n)_{k}\Bigg{\{}\sum_{\omega\in K_{\alpha}\setminus\{1\}}(x\omega)^{k}(1-x+x\omega)^{\alpha n-k}\Bigg{\}}$
$\hskip 28.45274pt-\Bigg{\{}\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha nx}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
$\hskip 28.45274pt+(-1)^{m}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha n(1-x)}^{(m)}(-t)}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t\Bigg{\}}$
$=:\mathrm{I}_{\alpha,m}^{(n)}(x)+\mathrm{II}_{\alpha,m}^{(n)}(x)-\mathrm{III}_{\alpha,m}^{(n)}(x).$ (2.31)
@@
$\mathrm{III}_{\alpha,m}^{(n)}(1-x)$ $=\frac{\alpha(1-x)^{\alpha}x^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha n(1-x)}^{(m)}(-t)}{|(tx)^{\alpha}-(1-x)^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
$\hskip 28.45274pt+(-1)^{m}\frac{\alpha(1-x)^{\alpha(n+1)}x^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha nx}^{(m)}(-t)}{|x^{\alpha}e^{-i\alpha\pi}-(t(1-x))^{\alpha}|^{2}}\,\mathrm{d}t$
$=(-1)^{m}\mathrm{III}_{\alpha,m}^{(n)}(x)$
for $x\in(0,1)$. Furthermore, $\mathrm{II}_{\alpha,m}(x)=0$ if $\alpha\in(0,2)$ because $K_{\alpha}=\{1\}$. Therefore, it follows from (2.31) that for $\alpha\in(0,2)$,
@@
$\mathrm{II}_{\alpha,m}^{(n)}(1-x)=(-1)^{m}\mathrm{II}_{\alpha,m}^{(n)}(x),\qquad\alpha>0.$ (2.33)
@@
Step 2. We give a uniform estimate of $\mathrm{II}_{\alpha,m}^{(n)}(x)$. Suppose that $\omega\in K_{\alpha}\setminus\{1\}$ and $x\in(0,1/2]$. For fixed $k\in\mathbb{N}_{0}$ with $0\leq k\leq m$, we have
@@
$\sup_{x\in(0,1/2]}|\mathrm{II}_{\alpha,m}^{(n)}(x)|\leq(m+1)(\#(K_{\alpha})-1)C_{2,5}\leq C_{2,6},$
@@
$\sup_{x\in(0,1)}|\mathrm{II}_{\alpha,m}^{(n)}(x)|=\sup_{x\in(0,1/2]}|\mathrm{II}_{\alpha,m}^{(n)}(x)|\leq C_{2,6}.$ (2.34)
Step 3. We next give a uniform estimate of $\mathrm{III}_{\alpha,m}^{(n)}(x)$. We write
$\mathrm{III}_{\alpha,m}^{(n)}(x)=\widetilde{\mathrm{III}}_{\alpha,m}^{(n)}(x)+(-1)^{m}\widetilde{\mathrm{III}}_{\alpha,m}^{(n)}(1-x)$
@@
$\widetilde{\mathrm{III}}_{\alpha,m}^{(n)}(x)=\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha nx}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t.$
@@
$|\widetilde{\mathrm{III}}_{\alpha,m}^{(n)}(x)|$ $\leq C_{2,8}\sum_{k=0}^{m}x^{\alpha}(1-x)^{\alpha(n+1)}|\sin\alpha\pi|$
×$\hskip 28.45274pt\times\{(nx)^{m-k}+1\}\int_{0}^{1}\frac{t^{\alpha-1}n^{k}t^{k}(1-t)^{\alpha n-k}}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t.$
@@
$|\widetilde{\mathrm{III}}_{\alpha,m}^{(n)}(x)|$ $\leq C_{2,10}\sum_{k=0}^{m}(1-x)^{\alpha n}\{(nx)^{m-k}+1\}$
$\leq C_{2,10}\sum_{k=0}^{m}\{(1-x)^{\alpha n}(nx)^{m-k}+(1-x)^{\alpha n}\}$
@@
$\sup_{x\in(0,1)}|\mathrm{III}_{\alpha,m}^{(n)}(x)|\leq C_{2,12},$ (2.35)
@@
$\sup_{x\in(0,1)}|(\alpha n)^{m}M_{\alpha,m}^{(n)}(x)-\mathrm{I}_{\alpha,m}^{(n)}(x)|=\sup_{x\in(0,1)}|\mathrm{II}_{\alpha,m}^{(n)}(x)-\mathrm{III}_{\alpha,m}^{(n)}(x)|\leq C_{2,13}.$
@@
We also have the following by applying Theorem 2.11 and (2.23).
@@
$\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}\left(\frac{j}{n}-x\right)^{m}\right|=\begin{cases}\displaystyle O\left(\frac{1}{n}\right)&\text{if }m=1,2,\vspace{2mm}\\ \displaystyle O\left(\frac{1}{n^{2}}\right)&\text{if }m=3,4.\end{cases}$
2.5. Characteristic function of the fractional binomial distribution
@@
×$\hskip 14.22636pt\times\Bigg{\{}\frac{1}{(t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi})(t^{\alpha}-\lambda^{\alpha}e^{i\alpha\pi})}+\frac{\lambda^{\alpha(n-2)}}{(\lambda^{-\alpha}e^{-i\alpha\pi}-t^{\alpha})(\lambda^{-\alpha}e^{i\alpha\pi}-t^{\alpha})}\Bigg{\}}\,\mathrm{d}t$ (2.37)
@@
We assume $\alpha\in\mathbb{N}$. Then, (2.36) holds for all $\lambda\in(0,1]$ by Proposition 1.2. Because both sides of (2.36) are holomorphic on $\mathbb{C}$, (2.36) holds for all $\lambda\in\mathbb{C}$ by analytic continuation.
@@
where $\arg(e^{\mp i\alpha\pi})$ is taken from $(0,2\pi)$. Because $\arg(e^{-i\alpha\pi})=(\overline{\alpha}-\alpha)\pi$ and $\arg(e^{i\alpha\pi})=(\alpha-\underline{\alpha})\pi$, (2.38) is rewritten as $-\eta_{\alpha}<\theta<\eta_{\alpha}$, where
$\eta_{\alpha}=-\big{[}\{2-(\alpha-\underline{\alpha})\}\wedge\{2-(\overline{\alpha}-\alpha)\}\wedge(\overline{\alpha}-\alpha)\wedge(\alpha-\underline{\alpha})\big{]}\frac{\pi}{\alpha}.$
Noting that $\eta_{\alpha}=\theta_{\alpha}/\alpha$ because $\overline{\alpha}-\underline{\alpha}=2$, both sides of (2.37) turn out to be holmorphic in the domain $D_{\alpha}$. Thus, (2.37) also holds for all $\lambda\in D_{\alpha}$ by analytic continuation. ∎
@@
$\geq 2b^{2}\sin^{2}\frac{\theta}{2}\geq(a^{2}+b^{2})\sin^{2}\frac{\theta}{2}.\qed$
@@
$\hskip 28.45274pt-\frac{1}{Z_{\alpha,x}^{(n)}}\frac{\alpha e^{i\xi}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
×$\hskip 28.45274pt\times\Bigg{[}\frac{x^{\alpha}(1-x)^{\alpha(n+1)}}{\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi-\alpha\pi)}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi+\alpha\pi)}\}}$
$\hskip 28.45274pt+\frac{x^{\alpha(n+1)}(1-x)^{\alpha}e^{i\xi(n-2)}}{\{(1-x)^{\alpha}e^{-i(\xi+\alpha\pi)}-(tx)^{\alpha}\}\{(1-x)^{\alpha}e^{-i(\xi-\alpha\pi)}-(tx)^{\alpha}\}}\Bigg{]}\,\mathrm{d}t$ (2.39)
@@
as $n\to\infty$ for some $\delta>0$ that is chosen uniformly on every compact subset of $x\in(0,1)$ and every compact subset of $\xi\in(-\theta_{\alpha},\theta_{\alpha})$. Here, if $\alpha\in\mathbb{N}$, then the last term of (2.39) is regarded as 00.
@@
×$\hskip 28.45274pt\times\left\{\frac{1}{(t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi})(t^{\alpha}-\lambda^{\alpha}e^{i\alpha\pi})}+\frac{\lambda^{\alpha(n-2)}}{(\lambda^{-\alpha}e^{-i\alpha\pi}-t^{\alpha})(\lambda^{-\alpha}e^{i\alpha\pi}-t^{\alpha})}\right\}\,\mathrm{d}t\Bigg{]}$
@@
$\hskip 28.45274pt-\frac{1}{Z_{\alpha,x}^{(n)}}\frac{\alpha x^{\alpha}(1-x)^{\alpha(n-1)}e^{i\xi}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
×$\hskip 28.45274pt\times\Bigg{[}\frac{(1-x)^{2\alpha}}{\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi-\alpha\pi)}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi+\alpha\pi)}\}}$
$\hskip 28.45274pt+\frac{x^{\alpha n}(1-x)^{-\alpha(n-2)}e^{i\xi n}}{\{(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}e^{i\xi}\}\{(1-x)^{\alpha}e^{i\alpha\pi}-(tx)^{\alpha}e^{i\xi}\}}\Bigg{]}\,\mathrm{d}t,$
@@
Note that $|\xi|<\theta_{\alpha}$ implies $\sin(\xi\pm\alpha\pi)/2\neq 0$. From these estimates and (2.13), we obtain (2.40). The case for $\alpha\in\mathbb{N}$ is simpler to prove. ∎
2.6. The fractional multinomial distribution

### v3

@@ 2.3 Uniform bound of $Z_{\alpha,x}^{(n)}$
If $\alpha\in(0,1)$, then [12, Theorem 1.2] implies that $Z_{\alpha,x}^{(n)}\leq 1$ for $n\in\mathbb{N}$ and $x\in[0,1]$. On the other hand, Equation (1.7) in [10, Theorem 1.3] leads to $Z_{\alpha,x}^{(n)}\leq\alpha$ for $n\in\mathbb{N}$ and $x\in[0,1]$ when $\alpha\geq 1$. Hence, the uniform upper bounds of (2.25) and (2.26) have been obtained.
We next discuss the lower bounds of $Z_{\alpha,x}^{(n)}$. Because it holds that $Z_{\alpha,x}^{(n)}=Z_{\alpha,1-x}^{(n)}$ and $Z_{\alpha,0}^{(n)}=\alpha$, it suffices to consider only the case where $x\in(0,1/2]$. It follows from [12, Proposition 2.6] that
@@
$\alpha\sum_{j=0}^{n}\binom{\alpha n}{\alpha j}x^{\alpha j}(1-x)^{\alpha(n-j)}+1\in\begin{cases}\left(0,(1-\alpha)(1-x)^{\alpha n}\right)\subset(0,1-\alpha)&\text{if }\alpha\in(0,1),\\ \left((1-\alpha)(1-x)^{\alpha n},0\right)&\text{if }\alpha\in(1,2).\end{cases}$
@@
$\geq\alpha\left(1-\frac{1}{2}\right)^{2\alpha nx}=\alpha\times 2^{-2\alpha nx},\qquad n\in\mathbb{N},\,x\in(0,1/2]$ (2.28)
@@
$Z_{\alpha,x}^{(n)}\geq\alpha\times 2^{-2\gamma},\qquad n\in\mathbb{N},\,x\in(0,1/2]\,\,\text{with $\alpha nx\leq\gamma$.}$
We now deal with the upper bound of $\sum_{\omega\in K_{\alpha}\setminus\{1\}}(1-x+x\omega)^{\alpha n}$ for $n\in\mathbb{N}$ and $x\in(0,1/2]$ with $\alpha nx>\gamma$. From (2.18), we have
@@
for $n\in\mathbb{N}$ and $x\in(0,1/2]$, where we used $(K_{\alpha})=\lfloor\alpha/2\rfloor+\lceil\alpha/2\rceil\leq\alpha+1$ in the fourth line and
@@
for $n\in\mathbb{N}$ and $x\in(0,1/2]$. ∎
2.4 Uniform estimates for moments of the fractional binomial distribution
@@
$\qquad\qquad-(\alpha n)^{-m}\sum_{k=0}^{m}W_{-\alpha nx}(m,k)(\alpha n)_{k}x^{k}\Bigg|=O\left(\frac{1}{n^{m}}\right)$ (2.30)
@@
$\quad+\sum_{k=0}^{m}W_{-\alpha nx}(m,k)(\alpha n)_{k}\Bigg\{\sum_{\omega\in K_{\alpha}\setminus\{1\}}(x\omega)^{k}(1-x+x\omega)^{\alpha n-k}\Bigg\}$
$\quad-\Bigg\{\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha nx}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
$\quad+(-1)^{m}\frac{\alpha x^{\alpha(n+1)}(1-x)^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha n(1-x)}^{(m)}(-t)}{|(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}|^{2}}\,\mathrm{d}t\Bigg\}$
$=:\mathrm{I}_{\alpha,m}^{(n)}(x)+\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)-\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x).$ (2.31)
@@
$\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha,m}^{(n)}(1-x)$ $=\frac{\alpha(1-x)^{\alpha}x^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha n(1-x)}^{(m)}(-t)}{|(tx)^{\alpha}-(1-x)^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t$
$\quad+(-1)^{m}\frac{\alpha(1-x)^{\alpha(n+1)}x^{\alpha}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha nx}^{(m)}(-t)}{|x^{\alpha}e^{-i\alpha\pi}-(t(1-x))^{\alpha}|^{2}}\,\mathrm{d}t$
$=(-1)^{m}\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)$
for $x\in(0,1)$. Furthermore, $\mathrm{I\hskip-1.00006ptI}_{\alpha,m}(x)=0$ if $\alpha\in(0,2)$ because $K_{\alpha}=\{1\}$. Therefore, it follows from (2.31) that for $\alpha\in(0,2)$,
@@
$\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(1-x)=(-1)^{m}\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x),\qquad\alpha>0.$ (2.33)
@@
Step 2. We give a uniform estimate of $\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)$. Suppose that $\omega\in K_{\alpha}\setminus\{1\}$ and $x\in(0,1/2]$. For fixed $k\in\mathbb{N}_{0}$ with $0\leq k\leq m$, we have
@@
$\sup_{x\in(0,1/2]}|\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)|\leq(m+1)(\#(K_{\alpha})-1)C_{2,5}\leq C_{2,6},$
@@
$\sup_{x\in(0,1)}|\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)|=\sup_{x\in(0,1/2]}|\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)|\leq C_{2,6}.$ (2.34)
Step 3. We next give a uniform estimate of $\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)$. We write
$\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)=\widetilde{\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}}_{\alpha,m}^{(n)}(x)+(-1)^{m}\widetilde{\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}}_{\alpha,m}^{(n)}(1-x)$
@@
$\widetilde{\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}}_{\alpha,m}^{(n)}(x)=\frac{\alpha x^{\alpha}(1-x)^{\alpha(n+1)}\sin\alpha\pi}{\pi}\int_{0}^{1}\frac{t^{\alpha-1}h_{-\alpha nx}^{(m)}(-t)}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t.$
@@
$|\widetilde{\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}}_{\alpha,m}^{(n)}(x)|$ $\leq C_{2,8}\sum_{k=0}^{m}x^{\alpha}(1-x)^{\alpha(n+1)}|\sin\alpha\pi|$
$\quad\times\{(nx)^{m-k}+1\}\int_{0}^{1}\frac{t^{\alpha-1}n^{k}t^{k}(1-t)^{\alpha n-k}}{|(t(1-x))^{\alpha}-x^{\alpha}e^{-i\alpha\pi}|^{2}}\,\mathrm{d}t.$
@@
$|\widetilde{\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}}_{\alpha,m}^{(n)}(x)|$ $\leq C_{2,10}\sum_{k=0}^{m}(1-x)^{\alpha n}\{(nx)^{m-k}+1\}$
$=C_{2,10}\sum_{k=0}^{m}\{(1-x)^{\alpha n}(nx)^{m-k}+(1-x)^{\alpha n}\}$
@@
$\sup_{x\in(0,1)}|\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)|\leq C_{2,12},$ (2.35)
@@
$\sup_{x\in(0,1)}|(\alpha n)^{m}M_{\alpha,m}^{(n)}(x)-\mathrm{I}_{\alpha,m}^{(n)}(x)|=\sup_{x\in(0,1)}|\mathrm{I\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)-\mathrm{I\hskip-1.00006ptI\hskip-1.00006ptI}_{\alpha,m}^{(n)}(x)|\leq C_{2,13}.$
@@
We also have the following by applying Theorems 2.9 and 2.11, and (2.23).
@@
$\sup_{\alpha\in K}\max_{x\in[0,1]}\left|\mathbb{E}\left[\left(\frac{S_{\alpha,x}^{(n)}}{n}-x\right)^{m}\right]\right|=\begin{cases}\displaystyle O\left(\frac{1}{n}\right)&\text{if }m=1,2,\\ \displaystyle O\left(\frac{1}{n^{2}}\right)&\text{if }m=3,4.\end{cases}$
2.5 Characteristic function of the fractional binomial distribution
@@
$\quad\times\Bigg\{\frac{1}{(t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi})(t^{\alpha}-\lambda^{\alpha}e^{i\alpha\pi})}+\frac{\lambda^{\alpha(n-2)}}{(\lambda^{-\alpha}e^{-i\alpha\pi}-t^{\alpha})(\lambda^{-\alpha}e^{i\alpha\pi}-t^{\alpha})}\Bigg\}\,\mathrm{d}t$ (2.37)
@@
We assume $\alpha\in\mathbb{N}$. Then, (2.36) holds for all $\lambda\in(0,1]$ by Proposition 1.1. Because both sides of (2.36) are holomorphic on $\mathbb{C}$, (2.36) holds for all $\lambda\in\mathbb{C}$ by the identity theorem for holomorphic functions.
@@
where $\arg(e^{\mp i\alpha\pi})$ is taken from $(0,2\pi)$. Because $\arg(e^{-i\alpha\pi})=(\overline{\alpha}-\alpha)\pi$ and $\arg(e^{i\alpha\pi})=(\alpha-\underline{\alpha})\pi$, (2.38) is rewritten as $\eta_{\alpha}<\theta<\eta_{\alpha}$, where
$\eta_{\alpha}=\big[\{2-(\alpha-\underline{\alpha})\}\wedge\{2-(\overline{\alpha}-\alpha)\}\wedge(\overline{\alpha}-\alpha)\wedge(\alpha-\underline{\alpha})\big]\frac{\pi}{\alpha}.$
By noting that $\eta_{\alpha}=\theta_{\alpha}/\alpha$ because $\overline{\alpha}-\underline{\alpha}=2$, both sides of (2.37) turn out to be holomorphic in the domain $D_{\alpha}$. Thus, (2.37) also holds for all $\lambda\in D_{\alpha}$ by the identity theorem. ∎
@@
$\geq 2b^{2}\sin^{2}\frac{\theta}{2}\geq(a^{2}+b^{2})\sin^{2}\frac{\theta}{2}.$
@@
$\quad-\frac{1}{Z_{\alpha,x}^{(n)}}\frac{\alpha e^{i\xi}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
$\quad\times\Bigg[\frac{x^{\alpha}(1-x)^{\alpha(n+1)}}{\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi-\alpha\pi)}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi+\alpha\pi)}\}}$
$\qquad+\frac{x^{\alpha(n+1)}(1-x)^{\alpha}e^{i\xi(n-2)}}{\{(1-x)^{\alpha}e^{-i(\xi+\alpha\pi)}-(tx)^{\alpha}\}\{(1-x)^{\alpha}e^{-i(\xi-\alpha\pi)}-(tx)^{\alpha}\}}\Bigg]\,\mathrm{d}t$ (2.39)
@@
as $n\to\infty$ for some $\delta>0$ that is chosen uniformly on any compact subset of $x\in(0,1)$ and any compact subset of $\xi\in(-\theta_{\alpha},\theta_{\alpha})$. Here, if $\alpha\in\mathbb{N}$, then the last term of (2.39) is regarded as 0.
@@
$\quad\times\left\{\frac{1}{(t^{\alpha}-\lambda^{\alpha}e^{-i\alpha\pi})(t^{\alpha}-\lambda^{\alpha}e^{i\alpha\pi})}+\frac{\lambda^{\alpha(n-2)}}{(\lambda^{-\alpha}e^{-i\alpha\pi}-t^{\alpha})(\lambda^{-\alpha}e^{i\alpha\pi}-t^{\alpha})}\right\}\,\mathrm{d}t\Bigg]$
@@
$\quad-\frac{1}{Z_{\alpha,x}^{(n)}}\frac{\alpha x^{\alpha}(1-x)^{\alpha(n-1)}e^{i\xi}\sin\alpha\pi}{\pi}\int_{0}^{1}t^{\alpha-1}(1-t)^{\alpha n}$
$\quad\times\Bigg[\frac{(1-x)^{2\alpha}}{\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi-\alpha\pi)}\}\{(t(1-x))^{\alpha}-x^{\alpha}e^{i(\xi+\alpha\pi)}\}}$
$\qquad+\frac{x^{\alpha n}(1-x)^{-\alpha(n-2)}e^{i\xi n}}{\{(1-x)^{\alpha}e^{-i\alpha\pi}-(tx)^{\alpha}e^{i\xi}\}\{(1-x)^{\alpha}e^{i\alpha\pi}-(tx)^{\alpha}e^{i\xi}\}}\Bigg]\,\mathrm{d}t,$
@@
Note that $|\xi|<\theta_{\alpha}$ implies $\sin(\xi\pm\alpha\pi)/2\neq 0$. From these estimates and Proposition 2.6 (1), we obtain (2.40). The case for $\alpha\in\mathbb{N}$ is simpler to prove. ∎
2.6 The fractional multinomial distribution